OpenStudy (anonymous):

If f(x) = x2 + 1, find f(3).

OpenStudy (amistre64):

what is 3^2 ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

f(3) = (3)^2 +1

OpenStudy (amistre64):

good, and 9+1?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

f(3) = 3^2 + 1 = 9 + 1 = 10

OpenStudy (amistre64):

then when x=3; f(3) = 10

OpenStudy (anonymous):

you just have to plug that number in to any x-value.

OpenStudy (anonymous):

Ohhhhhhhh! I'm such a duck.

OpenStudy (anonymous):

10 all you have to do is plug 3 into the (x) and then solve the problem out

OpenStudy (anonymous):

Be happy that you have so many persons helping you out, Nulify!